# Stationary Policies are Optimal in Risk-averse Total-reward MDPs with EVaR

### v1

@@ D.2 Bellman Operator
The exponential Bellman operator is monotone. That is for $\bm{x},\bm{y}\in\mathbb{R}^{S}$
@@
The property in (22) follows immediately from non-negattivity of $\bm{B}^{\bm{d}}$. The property in (23) then follows from the monotonicity of the $\max$ operator. ∎
@@
The lemma follows directly from the continuity of linear operators and from the fact that the pointwise maximum of a finite number of continuous functions is continuous. See also (Patek 2001, lemma 5)
@@
It is easy to see that $\bm{d}$ can be chosen independently for each state to maximize $\bm{v}$ uniformly across states. The optimality of deterministic decision rules, $\bm{d}\in\operatorname{ext}\mathcal{D}$, follows because ERM is a mixture quasi-convex function (Delage, Kuhn, and Wiesemann 2019).
The existence of value function for finite-horizon problem under the ERM objective has been analyzed previously (Hau, Petrik, and Ghavamzadeh 2023) including in the context of exponential utility functions (Chung and Sobel 1987).
To derive the exponential Bellman operator for the exponential value function for $\bm{d}\colon\mathcal{S}\to\Delta_{A}$ we concatenate the Bellman operator with the transformations to and from the exponential value function:
@@
[…]
@@
The existence of an optimal $\pi^{\star}$ follows by choosing the maximum in the definition of $L^{\star}$ which is attained by compactness and continuity of the objective. ∎
@@
$\bm{w}^{\infty}(\pi)=\bm{w}^{\infty}(\pi,\bm{z})=L^{\bm{d}}\bm{w}>-\bm{\infty}.$
@@
The result follows by algebraic manipulation from (12) and basic matrix analysis. When $\rho(\bm{B}^{\bm{d}})<1$, we get from Neumann series (Horn and Johnson 2013, problem 5.6.P26)
$\sum_{k=0}^{t-1}(\bm{B}^{\bm{d}})^{k}\bm{b}^{\bm{d}}=(\bm{I}-\bm{B}^{\bm{d}})^{-1}\bm{b}^{\bm{d}}$
and a consequence of Gelfand’s formula (Kallenberg 2021, theorem 4.5)
@@
When $\bm{\mu}^{\top}\bm{v}^{\infty,\star}=-\infty$ then the result follows immediately because $-\infty=\max_{\pi\in\Pi_{\mathrm{MR}}}\bm{\mu}^{\top}\bm{v}^{\infty}(\pi)\geq\max_{\pi\in\Pi_{\mathrm{SR}}}\bm{\mu}^{\top}\bm{v}^{\infty}(\pi)\geq-\infty$.
For the remainder of the proof, recall that $\pi_{\mathrm{M}}^{\star}\in\arg\max_{\pi\in\Pi_{\mathrm{MR}}}\bm{\mu}^{\top}\bm{v}^{\infty}(\pi)$ and suppose that $\bm{\mu}^{\top}\bm{v}^{\infty}(\pi^{\star}_{\mathrm{M}})>-\infty$. Then, the exponential value function $\bm{w}^{t,\star}=\bm{w}^{t}(\pi_{\mathrm{M}}^{\star})\in\mathbb{R}^{S}$ of $\pi_{\mathrm{M}}^{\star}$ satisfies by Theorem 3.1 that
@@
where the inequality follows from the inductive assumption and Lemma D.3.
Then, using the Monotone Convergence Theorem (Johnsonbaugh and Pfaffenberger 1981, theorem 16.2), finite $\mathcal{S}$, and $\inf_{t\in\mathbb{N}}\bm{w}_{\textrm{u}}^{t}\geq\inf_{t\in\mathbb{N}}\bm{w}^{t,\star}>-\infty$, we get that there exists $\bm{w}_{\mathrm{u}}^{\star}\in\mathbb{R}^{S}$ such that
@@
$\bm{w}_{\mathrm{u}}^{t}\;\geq\;\bm{w}^{t,\star}\;\geq\;\bm{w}_{\mathrm{l}}^{t},\qquad t\in\mathbb{N}.$
The Squeeze Theorem (Johnsonbaugh and Pfaffenberger 1981, theorem 14.3) then shows that
@@
We use $\bm{p}^{\bm{d}}\in\mathbb{R}_{+}^{S}$ to represent the probability of terminating in any state for each $\bm{d}\in\mathcal{D}$:
@@
Assume a policy $\pi=(\bm{d})_{\infty}\in\Pi_{\mathrm{SR}}$. Then, the probability of termination in $t\in\mathbb{N}$ or fewer steps is
$\sum_{k=0}^{t}\bm{\mu}^{\top}(\bm{P}^{\bm{d}})^{k}\bm{p}^{\bm{d}}\;=\;\bm{\mu}^{\top}(\bm{I}-(\bm{P}^{\bm{d}})^{t+1})\bm{1}.$
@@
$\bm{p}^{\bm{d}}=(\bm{I}-\bm{P})\bm{1}.$
The probability of terminating in step $t\geq 0$ exactly is
$\bm{\mu}^{\top}(\bm{P}^{\bm{d}})^{t}\bm{p}^{\bm{d}}.$
@@
Assume a $\pi=(\bm{d})_{\infty}\in\Pi_{\mathrm{SR}}$. Then there exists an eigenvector $\bm{f}\in\mathbb{R}^{S}$ such that $\bm{f}^{\top}\bm{B}^{\bm{d}}=\rho(\bm{B}^{\bm{d}})\cdot\bm{f}^{\top}$ and $\bm{f}\geq\bm{0},\bm{f}\neq\bm{0}$ and
$\bm{f}^{\top}\bm{b}^{\bm{d}}>0.$
@@
Because $\bm{B}^{\bm{d}}$ is non-negative, the required vector $\bm{f}$ exists from the Perron-Frobenius theorem, e.g. (Horn and Johnson 2013, Theorem 8.3.1). Therefore,
$\bm{f}^{\top}\bm{b}^{\bm{d}}\geq\bm{0},$
since $\bm{b}^{\bm{d}}\geq\bm{0}$.
It remains to show that $\bm{f}^{\top}\bm{b}^{\bm{d}}\neq 0$, which we do by deriving a contradiction. Without loss of generality, assume that $\bm{1}^{\top}\bm{f}=1$ and suppose that $\bm{f}^{\top}\bm{b}^{\bm{d}}=0$. Then:
@@
[…]
which is a contradiction with $\bm{1}^{\top}\bm{f}\neq 0$. The last step in the derivation follows from $\rho(\bm{P}^{\bm{d}})<1$ and therefore $\lim_{t\to\infty}(\bm{P}^{\bm{d}})^{t+1}=\bm{0}$ (Kallenberg 2021, Theorem 4.5). ∎
@@ E.1 Helper Lemmas
This well-known result follows, for example, from the dual representation of ERM (Hau, Petrik, and Ghavamzadeh 2023). ∎
@@
The proof uses the same technique as the lemma D.6 from (Hau, Petrik, and Ghavamzadeh 2023). We restate the proof here for completeness.
Using the fact that $f$ is non-increasing and $\beta\mapsto b^{-1}\log\alpha$ is increasing we get that:
$\sup_{\beta\in[\beta_{k},\beta_{k+1})}\left(f(\beta)+\beta^{-1}\cdot\log(\alpha)\right)-\left(f(\beta_{k})+\beta_{k}^{-1}\cdot\log\alpha\right)$ $\leq\sup_{\beta\in[\beta_{k},\beta_{k+1})}\left(f(\beta)+\beta_{k+1}^{-1}\cdot\log\alpha\right)-\left(f(\beta_{k})+\beta_{k}^{-1}\cdot\log\alpha\right)$
@@
The proof uses the same technique as the lemma D.6 from (Hau, Petrik, and Ghavamzadeh 2023). We restate the proof here for completeness.
@@
$\sup_{\beta\in[\beta_{K},\infty)}(f(\beta)+\beta^{-1}\log\alpha)-(f(\beta_{K})+\beta_{K}^{-1}\log\alpha)$ $\leq sup_{\beta\in[\beta_{K},\infty)}(f(\beta))-(f(\beta_{K})+\beta_{K}^{-1}\log\alpha)$ $\leq\frac{-\log(\alpha)}{\beta_{K}}.$
@@
$\max_{\beta\in\mathcal{B}(\beta_{0},\delta)}f(\beta)+\beta^{-1}\log\alpha\;\leq\;\sup_{\beta>0}f(\beta)+\beta^{-1}\log\alpha\;\leq\;\max_{\beta\in\mathcal{B}(\beta_{0},\delta)}f(\beta)+\beta^{-1}\log\alpha+\max\left\{f(0)-f(\beta_{0}),\delta\right\}.$
@@
E.2 Proof of Theorem 4.1
@@
$\sup_{\pi\in\Pi_{\mathrm{HR}}}\lim_{t\to\infty}\sup_{\beta>0}h_{t}(\pi,\beta)$ $\leq\lim_{t\to\infty}\sup_{\beta>0}\sup_{\pi\in\Pi_{\mathrm{HR}}}h_{t}(\pi,\beta)\stackrel{{\scriptstyle\text{(a)}}}{{\leq}}\lim_{t\to\infty}\sup_{\beta>0}h_{t}^{\star}(\beta)$
$\stackrel{{\scriptstyle\text{(b)}}}{{\leq}}\lim_{t\to\infty}\max_{\beta\in\mathcal{B}(\beta_{0},\delta)}h_{t}^{\star}(\beta)+\delta\stackrel{{\scriptstyle\text{(c)}}}{{\leq}}\max_{\beta\in\mathcal{B}(\beta_{0},\delta)}\lim_{t\to\infty}h_{t}^{\star}(\beta)+\delta$
@@
Here, (a) follows from Theorem 3.3, (b) from Lemma E.6, (c) from the continuity of $\max$ over finite sets.
@@
Then continuing the upper bound on the objective:
$\max_{\beta\in\mathcal{B}(\beta_{0},\delta)}h_{\infty}^{\star}(\beta)+\delta$ $=h_{\infty}(\pi^{\star},\beta^{\star})+\delta=\lim_{t\to\infty}h_{t}(\pi^{\star},\beta^{\star})+\delta$
$\leq\lim_{t\to\infty}\sup_{\beta>0}h_{t}(\pi^{\star},\beta)+\delta\leq\lim_{t\to\infty}\operatorname{EVaR}_{\alpha}^{\pi^{\star},\bm{\mu}}\left[\sum_{k=0}^{t}r(\tilde{s}_{k},\tilde{a}_{k},\tilde{s}_{k+1})\right]+\delta,$
which shows that $\pi^{\star}$ is $\delta$-suboptimal since $\lim_{t\to\infty}\operatorname{EVaR}_{\alpha}^{\pi^{\star},\bm{\mu}}\left[\sum_{k=0}^{t}r(\tilde{s}_{k},\tilde{a}_{k},\tilde{s}_{k+1})\right]$ is a lower bound on the objective. ∎
E.3 Proof of Corollary 4.2

### v2

@@ D.2 Bellman Operator
The exponential Bellman operator is monotone. That is, for $\bm{x},\bm{y}\in\mathbb{R}^{S}$
@@
The property in (24) follows immediately from non-negativity of $\bm{B}^{\bm{d}}$. The property in (25) then follows from the monotonicity of the $\max$ operator. ∎
@@
The lemma follows directly from the continuity of linear operators and from the fact that the pointwise maximum of a finite number of continuous functions is continuous. See also (Patek, 2001, lemma 5)
@@
It is easy to see that $\bm{d}$ can be chosen independently for each state to maximize $\bm{v}$ uniformly across states. The optimality of deterministic decision rules, $\bm{d}\in\operatorname{ext}\mathcal{D}$, follows because ERM is a mixture quasi-convex function Delage et al. (2019).
The existence of value function for finite-horizon problem under the ERM objective has been analyzed previously Hau et al. (2023b) including in the context of exponential utility functions Chung and Sobel (1987).
To derive the exponential Bellman operator for the exponential value function for $\bm{d}\colon\mathcal{S}\to\Delta_{A}$, we concatenate the Bellman operator with the transformations to and from the exponential value function:
@@
$\;=\;-\mathbb{E}^{\bm{d},s}\left[\operatorname{exp}\left(-\beta\cdot r(s,\tilde{a}_{0},\tilde{s}_{1})+\log(-w_{\tilde{s}_{1}})\right)\right]$
$\;=\;-\mathbb{E}^{\bm{d},s}\left[\operatorname{exp}\left(-\beta\cdot r(s,\tilde{a}_{0},\tilde{s}_{1})\right)\cdot(-w_{\tilde{s}_{1}})\right]$
$\;=\;\sum_{s^{\prime}\in\bar{\mathcal{S}}}\sum_{a\in\mathcal{A}}p(s,a,s^{\prime})\cdot d_{a}(s)\cdot\operatorname{exp}\left(-\beta\cdot r(s,a,s^{\prime})\right)\cdot w_{s^{\prime}}$
$\;=\;\sum_{s^{\prime}\in\mathcal{S}}\sum_{a\in\mathcal{A}}p(s,a,s^{\prime})\cdot d_{a}(s)\cdot\operatorname{exp}\left(-\beta\cdot r(s,a,s^{\prime})\right)\cdot w_{s^{\prime}}$
@@
The existence of an optimal $\pi^{\star}$ follows by choosing the maximum in the definition of $L^{\star}$, which is attained by compactness and continuity of the objective. ∎
@@
$\bm{w}^{\infty}(\pi)=\bm{w}^{\infty}(\pi,\bm{z})=L^{\bm{d}}\bm{w}^{\infty}(\pi)>-\bm{\infty}.$
@@
The result follows by algebraic manipulation from (12) and basic matrix analysis. When $\rho(\bm{B}^{\bm{d}})<1$, we get from Neumann series (Horn and Johnson, 2013, problem 5.6.P26)
$\lim_{t\to\infty}\sum_{k=0}^{t-1}(\bm{B}^{\bm{d}})^{k}\bm{b}^{\bm{d}}=(\bm{I}-\bm{B}^{\bm{d}})^{-1}\bm{b}^{\bm{d}},$
and a consequence of Gelfand’s formula (Kallenberg, 2021, theorem 4.5)
@@
For the remainder of the proof, let $\pi_{\mathrm{M}}^{\star}\in\arg\max_{\pi\in\Pi_{\mathrm{MR}}}\bm{\mu}^{\top}\bm{v}^{\infty}(\pi)$ and suppose that $\bm{\mu}^{\top}\bm{v}^{\infty}(\pi^{\star}_{\mathrm{M}})>-\infty$. Then, the exponential value function $\bm{w}^{t,\star}=\bm{w}^{t}(\pi_{\mathrm{M}}^{\star})\in\mathbb{R}^{S}$ of $\pi_{\mathrm{M}}^{\star}$ satisfies by Theorem 3.1 that
@@
where the inequality follows from the inductive assumption and from Lemma D.3.
Then, using the Monotone Convergence Theorem (Johnsonbaugh and Pfaffenberger, 1981, theorem 16.2), finite $\mathcal{S}$, and $\inf_{t\in\mathbb{N}}\bm{w}_{\textrm{u}}^{t}\geq\inf_{t\in\mathbb{N}}\bm{w}^{t,\star}>-\infty$, we get that there exists $\bm{w}_{\mathrm{u}}^{\star}\in\mathbb{R}^{S}$ such that
@@
$\bm{w}_{\mathrm{u}}^{t}\;\geq\;\bm{w}^{t,\star}\;\geq\;\bm{w}_{\mathrm{l}}^{t},\qquad t\in\mathbb{N},$
the Squeeze Theorem (Johnsonbaugh and Pfaffenberger, 1981, theorem 14.3) shows that
@@
We use $\bm{p}^{\bm{d}}\in\mathbb{R}_{+}^{S}$ to represent the probability of terminating from any state for each $\bm{d}\in\mathcal{D}$:
@@
Assume a policy $\pi=(\bm{d})_{\infty}\in\Pi_{\mathrm{SR}}$. Then, the probability of terminating in $t\in\mathbb{N},t>0$ or fewer steps is
$\sum_{k=0}^{t-1}\bm{\mu}^{\top}(\bm{P}^{\bm{d}})^{k}\bm{p}^{\bm{d}}\;=\;\bm{\mu}^{\top}(\bm{I}-(\bm{P}^{\bm{d}})^{t})\bm{1},$
where $\bm{P}^{\bm{d}}$ is a $|\mathcal{S}|\times|\mathcal{S}|$ matrix, and each element $P^{\bm{d}}_{s,s^{\prime}}=\sum_{a\in\mathcal{A}}d_{a}(s)\cdot p(s,a,s^{\prime}),\forall s,s^{\prime}\in\mathcal{S}$.
@@
$\bm{p}^{\bm{d}}=(\bm{I}-\bm{P}^{\bm{d}})\bm{1}.$
The probability of terminating in step $t>0$ exactly is
$\bm{\mu}^{\top}(\bm{P}^{\bm{d}})^{t-1}\bm{p}^{\bm{d}}.$
@@
For each $\pi=(\bm{d})_{\infty}\in\Pi_{\mathrm{SR}}$, there exists an $\bm{f}\in\mathbb{R}^{S}$ such that
$\bm{f}^{\top}\bm{B}^{\bm{d}}=\rho(\bm{B}^{\bm{d}})\cdot\bm{f}^{\top},\quad\bm{f}\geq\bm{0},\quad\bm{f}\neq\bm{0},\quad\bm{f}^{\top}\bm{b}^{\bm{d}}>0.$
@@
Because $\bm{B}^{\bm{d}}$ is non-negative, there exists an $\bm{f}\in\mathbb{R}^{S}$ from the Perron-Frobenius theorem (Horn and Johnson, 2013, Theorem 8.3.1), such that
$\bm{f}^{\top}\bm{B}^{\bm{d}}=\rho(\bm{B}^{\bm{d}})\cdot\bm{f}^{\top},\qquad\bm{f}\geq\bm{0},\qquad\bm{f}\neq\bm{0}.$
Furthermore, because $\bm{b}^{\bm{d}}\geq\bm{0}$, we also have that
$\bm{f}^{\top}\bm{b}^{\bm{d}}\geq\bm{0}.$
To prove the theorem, it remains to show that $\bm{f}^{\top}\bm{b}^{\bm{d}}\neq 0$, which we do by deriving a contradiction. Note that the $\bm{f}$ constructed from the Perron-Frobenius theorem is scale-independent. Therefore, without loss of generality, assume that $\bm{1}^{\top}\bm{f}=1$ and suppose that $\bm{f}^{\top}\bm{b}^{\bm{d}}=0$. Then:
@@
$\bm{f}^{\top}(\bm{P}^{\bm{d}})^{k}\bm{p}^{\bm{d}}$ $=0,\,\forall k\in\mathbb{N}$ [ from Lemma D.8 ]
$\sum_{k=0}^{t-1}\bm{f}^{\top}(\bm{P}^{\bm{d}})^{k}\bm{p}^{\bm{d}}$ $=0,\,\forall t\in\mathbb{N},t>0$ [ by summing elements ]
$\bm{f}^{\top}(\bm{I}-(\bm{P}^{\bm{d}})^{t})\bm{1}$ $=0,\,\forall t\in\mathbb{N}$ [ from Lemma D.6 ]
$\lim_{t\to\infty}\bm{f}^{\top}(\bm{I}-(\bm{P}^{\bm{d}})^{t})\bm{1}$ $=0,$ [ limit ]
$\bm{f}^{\top}\bm{1}$ $=0,$ [ from Lemma 2.3 ]
which is a contradiction with $\bm{1}^{\top}\bm{f}\neq 0$. The last step in the derivation follows from $\rho(\bm{P}^{\bm{d}})<1$ and therefore $\lim_{t\to\infty}(\bm{P}^{\bm{d}})^{t}=\bm{0}$ (Kallenberg, 2021, Theorem 4.5). ∎
@@ E.1 Helper Lemmas
This well-known result follows, for example, from the dual representation of ERM Hau et al. (2023b). ∎
@@
The proof uses the same technique as the lemma D.5 from Hau et al. (2023b). We restate the proof here for completeness.
Using the fact that $f$ is non-increasing and $\beta\mapsto\beta^{-1}\log\alpha$ is increasing we get that:
$\sup_{\beta\in[\beta_{k},\beta_{k+1})}\left(f(\beta)+\beta^{-1}\cdot\log(\alpha)\right)-\left(f(\beta_{k})+\beta_{k}^{-1}\cdot\log\alpha\right)$
$\leq\sup_{\beta\in[\beta_{k},\beta_{k+1})}\left(f(\beta)+\beta_{k+1}^{-1}\cdot\log\alpha\right)-\left(f(\beta_{k})+\beta_{k}^{-1}\cdot\log\alpha\right)$
@@
The proof uses the same technique as the lemma D.6 from Hau et al. (2023b). We restate the proof here for completeness.
@@
$\sup_{\beta\in[\beta_{K},\infty)}(f(\beta)+\beta^{-1}\log\alpha)-(f(\beta_{K})+\beta_{K}^{-1}\log\alpha)\leq\sup_{\beta\in[\beta_{K},\infty)}(f(\beta))-(f(\beta_{K})+\beta_{K}^{-1}\log\alpha)\leq\frac{-\log(\alpha)}{\beta_{K}}.$
@@
$\max_{\beta\in\mathcal{B}(\beta_{0},\delta)}f(\beta)+\beta^{-1}\log\alpha$ $\;\leq\;\sup_{\beta>0}f(\beta)+\beta^{-1}\log\alpha$
$\;\leq\;\max_{\beta\in\mathcal{B}(\beta_{0},\delta)}f(\beta)+\beta^{-1}\log\alpha+\max\left\{f(0)-f(\beta_{0}),\delta\right\}.$
@@
E.2 Proof of Proposition 4.1
Proof of Proposition 4.1.
First, the iterates $\beta_{k},k\geq 0$ can be bounded as
$\beta_{k+1}$ $=\frac{\beta_{k}\log\frac{1}{\alpha}}{\log\frac{1}{\alpha}-\beta_{k}\delta}\geq\beta_{k}\frac{\log\frac{1}{\alpha}}{\log\frac{1}{\alpha}-\beta_{0}\delta}$
$=\beta_{0}\left(\frac{\log\frac{1}{\alpha}}{\log\frac{1}{\alpha}-\beta_{0}\delta}\right)^{k+1}.$
Using the lower bound on $\beta_{k}$, to show that
$\beta_{K}\geq\frac{\log\frac{1}{\alpha}}{\delta},$
it is sufficient to show that
$\beta_{0}\left(\frac{\log\frac{1}{\alpha}}{\log\frac{1}{\alpha}-\beta_{0}\delta}\right)^{K}\geq\frac{\log\frac{1}{\alpha}}{\delta}.$
Using the variable $z$, the sufficient condition translates to
$\left(\frac{1}{1-z}\right)^{K}$ $\geq\frac{1}{z}$
$K$ $\geq\frac{\log z}{\log(1-z)},$
since $z\in(0,1)$ by the assumption $\beta_{0}\delta<\log\frac{1}{\alpha}$. ∎
E.3 Proof of Theorem 4.2
@@
$\sup_{\pi\in\Pi_{\mathrm{HR}}}\lim_{t\to\infty}\sup_{\beta>0}h_{t}(\pi,\beta)$ $\stackrel{{\scriptstyle\text{(a)}}}{{\leq}}\lim_{t\to\infty}\sup_{\beta>0}\sup_{\pi\in\Pi_{\mathrm{HR}}}h_{t}(\pi,\beta)$
=$\stackrel{{\scriptstyle\text{(b)}}}{{=}}\lim_{t\to\infty}\sup_{\beta>0}h_{t}^{\star}(\beta)$
$\stackrel{{\scriptstyle\text{(c)}}}{{\leq}}\lim_{t\to\infty}\max_{\beta\in\mathcal{B}(\beta_{0},\delta)}h_{t}^{\star}(\beta)+\delta$
$\stackrel{{\scriptstyle\text{(d)}}}{{\leq}}\max_{\beta\in\mathcal{B}(\beta_{0},\delta)}\lim_{t\to\infty}h_{t}^{\star}(\beta)+\delta$
@@
Here, (a) follows because the limit of an upper bound on a sequence is an upper bound on the limit, (b) follows from the definition of $h_{t}^{\star}$, (c) from Lemma E.6, (d) from the continuity of $\max$ over finite sets.
@@
Then, continuing the upper bound on the objective:
$\max_{\beta\in\mathcal{B}(\beta_{0},\delta)}h_{\infty}^{\star}(\beta)+\delta$ $=h_{\infty}(\pi^{\star},\beta^{\star})+\delta$
$=\lim_{t\to\infty}h_{t}(\pi^{\star},\beta^{\star})+\delta$
$\leq\lim_{t\to\infty}\sup_{\beta>0}h_{t}(\pi^{\star},\beta)+\delta$
$\leq\lim_{t\to\infty}\operatorname{EVaR}_{\alpha}^{\pi^{\star},\bm{\mu}}\left[\sum_{k=0}^{t-1}r(\tilde{s}_{k},\tilde{a}_{k},\tilde{s}_{k+1})\right]+\delta,$
which shows that $\pi^{\star}$ is $\delta$-suboptimal since $\lim_{t\to\infty}\operatorname{EVaR}_{\alpha}^{\pi^{\star},\bm{\mu}}\left[\sum_{k=0}^{t-1}r(\tilde{s}_{k},\tilde{a}_{k},\tilde{s}_{k+1})\right]$ is a lower bound on the objective. ∎
E.4 Proof of Corollary 4.3
@@
F Additional Material of Section 5
The detailed description of the gambler’s ruin is as follows. The state space is $\bar{\mathcal{S}}=\{0,1,\cdots,7,e\}$ and the action space is $\mathcal{A}=\{0,1,\cdots,6\}$. Action 00 represents quitting the game, and the other actions represent the bet size. The initial state is chosen randomly according to a uniform distribution over states $1,\dots,7$.
The transitions and rewards in the MDP are defined as follows.
In state 00, the gambler has capital 00, takes the only available action 00, receives the reward $-1$, and transits to the sink state $e$. That is:
$\bar{p}(0,0,e)=1,\qquad\bar{r}(0,0,e)=-1.$
If state $7$, the gambler has capital 7, reaches the target wealth level, takes the only available action 00, receives the reward $7$, and transits to the sink state $e$. That is:
$\bar{p}(7,0,e)=1,\qquad\bar{r}(7,0,e)=7.$
For the sink state $e$, the only available action is 00, and the transition probabilities and rewards are
$\bar{p}(e,0,e)=1,\qquad\bar{r}(e,0,e)=0.$
For other states $s\in\left\{1,\dots,6\right\}$, the gambler’s capital is $s$, can take any action $a\in\{1,\cdots,s\}$ to continue the game or an action 00 to quit the game.
If the game is won, the gambler’s bet doubles. If the game is lost, the gambler’s bet is lost. That is, the transition probabilities and rewards for $a\in\left\{1,\dots,s\right\}$ are
$\bar{p}(s,a,s^{\prime})=\begin{cases}q&\text{if }s^{\prime}=\min\left\{s+a,7\right\},\\ 1-q&\text{if }s^{\prime}=s-a,\\ 0&\text{otherwise},\end{cases}\qquad\bar{r}(s,a,s^{\prime})=0,$
for each $s^{\prime}\in\left\{0,\dots,7,e\right\}$ where $q\in[0,1]$ is the win probability.
If $s\in\left\{1,\dots,6\right\}$ and the action is 00, the gambler quits the game and collects the reward. That is:
$\bar{p}(s,0,e)=1,\qquad\bar{r}(s,0,e)=s.$